7. An aeroplane heads due north at 500 km/h. It experiences a 80 km/h crosswind flowing in


the direction N60oE.


(a) Find the true velocity of the aeroplane. (7)


(b) Determine the speed of the aeroplane. (Leave your answer in terms of square root


1
Expert's answer
2022-05-17T23:13:10-0400

"a^2=b^2+c^2-2bc cos \alpha"

"a^2=500^2+80^2-2(500)80cos(60\u00b0)=250000+6400-40000=216400"

true velocity of the aeroplane =speed of the aeroplane

"a=\sqrt{216400}"